Two consecutive angles of a parallelogram are in the ratio 43:47. How many degrees are there in each angle of the parallelogram?
Answer by ikleyn(52781)

.

So the angles are 43x and 47x, giving the sum of  43x + 47x = 90x,

where x is some, now unknown, common measure of the angles.


The sum of two consecutive angles in any parallelogram is 180 degrees, so

    90x = 180,  giving  x = 180/90 = 2 degrees.


Thus the angles are 2*43 = 86  degrees  and  2*47 = 94 degrees.    ANSWER

Solved.